OpenStudy (anonymous):

Will someone explain how to factor a trinomial? My practice work says: factor each trinomial. 3x^2-20x+28

OpenStudy (kirbykirby):

You multiply the 3 and 28 together: 3*28 = 84 Now you find all factors of 84, such that adding both of them will give -20 So, -6 and -14 work because (-6*-14)= 84 and -6 + (-14) = -20 So, what you do is write down 3x^2 -6x -14x + 28 [because -6x-14x = -20x, you break it down using the factors 06, -14 you found) Now, you factor the first 2 terms, and the last two terms: 3x(x-2) - 14(x-2); now you have (x-2) is a common factor, so you factor again: (x-2)(3x-14)

OpenStudy (anonymous):

That is correct! Thank you so much for your explanation!

OpenStudy (kirbykirby):

yw :)